OpenStudy (anonymous):

During season 7 of The Biggest Loser, Tara lost 15 lb more than Helen. The amount of weight that Mike lost was 73 lb less than twice Helen's weight loss. They lost a combined 502 lb. How much weight did each contestant lose?

OpenStudy (imstuck):

it seems that everyone's weight is based on helen's weight, so we will call her weight x. Tara lost 15 more than Helen, so Tara's weight will be labeled x + 15. Mike's weight loss is 73 less than twice Helen's, so Mike's loss is 2x - 73. Added all together it is 502 pounds. So the expressions added all together form the equation: (x) + (x + 15) + (2x - 73) = 502

OpenStudy (texaschic101):

T + H + M = 502 T = H + 15 M = 2H - 73 Now we sub (H + 15) + H + (2H - 73) = 502 can you take it from here ?

OpenStudy (anonymous):

shouldn't total should me equal 502 but i got 362

OpenStudy (anonymous):

oki got it